# Topological Photon Transport in Programmable Photonic Processors via Discretized Evolution of Synthetic Magnetic Fields

### v2

@@ Abstract
Photons, unlike electrons, do not couple directly to magnetic fields, yet synthetic gauge fields can impart magnetic-like responses and enable topological transport. Discretized Floquet evolution provides a controlled route, where the time-ordered sequencing of non-commuting Hamiltonians imprints complex hopping phases and breaks time-reversal symmetry. However, stabilizing such driven dynamics and observing unambiguous topological signatures on a reconfigurable platform has remained challenging. Here we demonstrate synthetic gauge fields for light on a programmable photonic processor by implementing discretized Floquet drives that combine static and dynamic phases. This approach reveals hallmark features of topological transport: chiral circulation that reverses under drive inversion, flux-controlled interference with high visibility, and robust directional flow stabilized by maximizing the minimal Floquet quasi-energy gap. The dynamics are further characterized by a first-harmonic phase order parameter, whose per-period winding number quantifies angular drift and reverses sign with the drive order. These results establish discretized, gap-optimized Floquet evolution as a versatile and fully programmable framework for topological photonics, providing a compact route to engineer gauge fields, stabilize driven phases, and probe winding-number signatures of chiral transport.
@@ 1. Introduction
Magnetic fields profoundly alter the behaviour of charged particles by reshaping their energy spectrum. In two-dimensional electron systems, a perpendicular field condenses continuous bands into a hierarchy of discrete Landau levels, each separated by the cyclotron frequency and hosting a macroscopic degeneracy of states. This quantization underlies the quantum Hall effect, wherein the electrical conductance assumes precisely quantized values. The effect also produces chiral edge states, which are one-way electron channels along material boundaries that are topologically protected against backscattering from disorder [1, 2]. Photons propagate through such magnetic fields unchanged, their spectra unquantized, their transport unprotected. Remarkably, synthetic magnetic fields can induce analogous phenomena in photonic systems, thereby enabling magnetic-like responses and robust optical transport [3, 4, 5].
@@
In this work, we demonstrate a photonic quantum simulator of synthetic magnetic fields based on discrete-time evolution. The simulator is implemented on a reconfigurable Mach-Zehnder interferometer (MZI) mesh, which supports arbitrary linear optical transformations [19, 20] and has enabled applications in quantum information processing [21, 22, 23, 24], neuromorphic computing [25, 26], and signal processing [27, 28]. We use a Floquet drive in which each period is divided into unitary substeps that sequentially activate selected mode pairs on the MZI mesh while other couplings are held off. The time-ordered product of these steps induces complex hopping phases for photons, realizing a synthetic magnetic flux. Because the substeps do not commute, the evolution depends on their order, which breaks time-reversal symmetry. This programmability allows a single photonic processor to investigate various topological phenomena without separate fabrications by simply adjusting control signals [29, 30]. Leveraging this versatility, we implement three-, four-, and seven-site lattice configurations. Our measurements reveal how each geometry produces distinct magnetic signatures, with the three-site system establishing the fundamental mechanism, the four-site system revealing flux interference, and the seven-site system demonstrating directional photon transport in complex systems.
@@ 2.1. Theoretical Framework
The generation of synthetic magnetic fields for photons in our programmable photonic processor relies on the principle of dynamically breaking time-reversal symmetry. This is achieved through a periodic, time-ordered sequence of hopping configurations. As this process constitutes a periodically driven system, its dynamics are described within the framework of Floquet theory.
This principle is best illustrated by analyzing a three-site triangular lattice, for which we define a set of hopping Hamiltonians $\{H_{1},H_{2},H_{3}\}$ that activate each edge sequentially:
$H_{1}=J(|1\rangle\langle 2|+|2\rangle\langle 1|),\quad H_{2}=J(|2\rangle\langle 3|+|3\rangle\langle 2|),\quad H_{3}=J(|3\rangle\langle 1|+|1\rangle\langle 3|),$ (1)
where $J$ is the coupling strength and $|i\rangle$ represents a photon localized at site $i$. These Hamiltonians cycle with period $T$, each operating for duration $T/3$ before switching to the next, creating a time-periodic system $H(t)=H(t+T)$. Each subprocess evolves the system through the unitary operator:
$U_{k}(T/3)=\exp(-iH_{k}T/3),\quad k=1,2,3.$ (2)
The Floquet operator, $U_{F}(T)$, is the time-ordered product of these operators. Consequently, the clockwise (CW) modulation, corresponding to the sequence $H_{1}\rightarrow H_{2}\rightarrow H_{3}$, is distinct from its reversed, counter-clockwise (CCW) analogue:
$U_{F}^{\text{CW}}(T)=U_{3}U_{2}U_{1},\qquad U_{F}^{\text{CCW}}(T)=U_{1}U_{2}U_{3}.$ (3)
For an arbitrary evolution time $t$, the total unitary operator $U(t)$ can be decomposed based on the number of full driving periods, $n$, and the residual time, $\tau$, such that $t=nT+\tau$ where $0\leq\tau<T$. The total evolution operator then takes the form:
$U(t)=U_{\text{partial}}(\tau)[U_{F}(T)]^{n},$ (4)
where $U_{F}(T)]^{n}$ represents the stroboscopic evolution, and $U_{\text{partial}}(\tau)$ accounts for the incomplete final period. This partial operator is the time-ordered product of the operators $U_{k}$ that fit within the duration $\tau$. The evolved state is then $|\psi(t)\rangle=U(t)|\psi(0)\rangle$, from the fixed initial state $|\psi(0)\rangle$. This discretized cyclic evolution through the Floquet operator is visualized in Fig. 1.
@@
Fig. 1: Discretized Floquet evolution on a programmable photonic processor. A state $|\psi(0)\rangle$ injected on the left bus enters a programmable photonic processor (black box) that implements the time-evolution operator $U(t)$ on NN modes, yielding $|\psi(t)\rangle=U(t)|\psi(0)\rangle$ at the right bus. a, Clockwise (CW) drive: one period $T$ is decomposed into three sequential substeps $U_{1}(T/3)$ (red), $U_{2}(T/3)$ (green), and $U_{3}(T/3)$ (blue). These time-ordered operations constitute the Floquet operator $U_{F}^{\text{CW}}(T)=U_{3}U_{2}U_{1}$. Dashed guide lines indicate that this schedule is programmed onto the chip; horizontal dots indicate repetition over successive periods. At intermediate times (e.g., $t=T/3$), only the substep executed up to that moment has acted. b, Counter-clockwise (CCW) drive obtained by reversing the order of the substeps, giving $U_{F}^{\text{CCW}}(T)=U_{1}U_{2}U_{3}$. c, Scalable NN-mode lattice representation. Solid bonds show the coupling pattern active in each substep (colors match panels a–b); dashed bond extensions indicate tiling to larger lattices; white circles mark the modes (waveguides). Curved arrows highlight the induced chiral circulation in each triangular plaquette, illustrating how the programmed drive order generates a synthetic magnetic flux.
The time-ordered Floquet drive induces a geometric phase for photons as they hop around the lattice, an effect analogous to the Aharonov-Bohm effect for charged particles in a magnetic field. To see how this emerges, we examine the stroboscopic evolution of the system which by Floquet’s theorem can be described by a time-independent effective Hamiltonian $H_{\text{eff}}$ defined through:
$U_{F}(T)=\exp(-iH_{\text{eff}}T)\quad\Rightarrow\quad H_{\text{eff}}=\frac{i}{T}\ln U_{F}(T).$ (5)
While the constituent Hamiltonians are real, the non-commuting drive sequence results in an effective Hamiltonian with complex hopping amplitudes:
$H_{\text{eff}}=\sum_{i}V_{\text{eff}}^{(i)}|i\rangle\langle i|+\sum_{\langle i,j\rangle}\left(J_{\text{eff}}^{(ij)}e^{i\phi_{ij}}|i\rangle\langle j|+\text{h.c.}\right).$ (6)
Here, $V_{\text{eff}}^{(i)}$ are emergent on-site potentials and $J_{\text{eff}}$ are effective coupling strengths. For larger lattices the effective Floquet Hamiltonian is not strictly short-range and, in principle, can include all-to-all terms; it is also not unique for a given $U_{F}(T)$. The phases $\phi_{ij}$ can be associated to a synthetic vector potential via:
$\phi_{ij}=\int_{\mathbf{r}_{i}}^{{\mathbf{r}_{j}}}\mathbf{A}\cdot d\mathbf{l},$ (7)
with the corresponding synthetic magnetic field $\mathbf{B}=\nabla\times\mathbf{A}.$
In addition to the dynamically programmed phase, $\Phi_{\text{mod}}=\sum\phi_{ij}$, which is set by the modulation timing, we can introduce statically programmed phases, $\Phi_{\text{static}}=\sum_{k}\alpha_{k}$, by modifying the hopping terms of the constituent Hamiltonians:
$H_{k}=J(e^{i\alpha_{k}}|i\rangle\langle j|+\text{h.c.}),$ (8)
so that both the dynamic and static contributions combine additively as the total Aharonov-Bohm phase accumulated in our lattice:
$\Phi_{\text{total}}=\Phi_{\text{mod}}+\Phi_{\text{static}}.$ (9)
The principle established here can be readily applied to engineer programmable synthetic magnetic fields in larger systems composed of interconnected triangular plaquettes (Fig. 1c).
@@ 2.2. Experimental Platform
A single programmable photonic chip allowed us to implement three distinct lattice configurations. We first realized a three-site triangular lattice whose single closed loop allowed clear interpretation of directional effects when driving CW and CCW, isolating the non-commuting operations that break time-reversal symmetry. We then programmed a four-site lattice consisting of two coupled triangular lattices with independent phase control. The shared vertices where pathways converge provided observation points for interference effects. Finally, we configured a seven-site hexagonal lattice to test synthetic magnetic fields in multi-loop geometries. Six perimeter sites surrounding a central node formed a network of interconnected triangular plaquettes, revealing whether directional transport persists in complex systems.
Operating conditions varied across lattice configurations based on their specific requirements. For coherent population transfer in the three-site and four-site systems, we set coupling strengths and evolution times to achieve Rabi $\pi$-pulses. This operating point is robust to perturbations in a broad range, given by a finite gap in the corresponding spectrum of the three-site Floquet operator, so the dynamics are tolerant to small period errors as shown in Fig. 5a. While the triangular lattice experiments tested both modulation sequences to establish chirality, the four-site system used only CW driving to focus solely on the interference effects. The seven-site system required a different regime where Rabi pulses are not applicable, so we selected the modulation period by maximizing the minimal quasi-energy gap of the Floquet operator, which yields a stable evolution and robust directional flow. As in the 3-site case, testing both modulation sequences would confirm the presence of the synthetic magnetic field signatures.
[…]
@@ 2.3. Unitary Implementation
We realize the evolution operators from Eqs. (3) by programming the MZI mesh on the photonic chip. Each MZI contains two 3-dB multimode interference couplers and two thermo-optic phase shifters, together implementing a two-mode SU(2) transformation with transfer matrix (Sec. S1):
$T_{\text{MZI}}=ie^{i\theta/2}\begin{bmatrix}e^{i\varphi}\sin(\theta/2)&\cos(\theta/2)\\ e^{i\varphi}\cos(\theta/2)&-\sin(\theta/2)\end{bmatrix}$, (10)
where the internal phase shift $\theta\in[0,\pi]$ controls the power splitting ratio between output ports as $\sin^{2}(\theta/2):\cos^{2}(\theta/2)$, and the external phase shift $\varphi\in[0,2\pi]$ sets their relative phase. Titanium nitride heaters on the upper waveguide arm generate these phase shifts through the thermo-optic effect, with the external heater positioned before the first coupler and the internal heater placed between the two couplers.
Since our theoretical models utilize only a subset of the 12 available modes, we embed the evolution operators into the full-dimensional space. For an $n$-site lattice with operator $U_{n}$ acting on selected modes, the embedded $12\times 12$ unitary takes the block-diagonal form:
$U_{12}=\begin{bmatrix}U_{n}&0\\ 0&I_{12-n}\end{bmatrix}$, (11)
where $I_{12-n}$ is the identity matrix for the unused modes. This ensures that light in inactive modes propagates unchanged while lattice dynamics occur exclusively in the active subspace. Physically, MZIs connecting to unused modes are set to maximum reflectivity $\theta=\pi$, preventing coupling between these modes.
The Clements decomposition [20] expresses this unitary as:
$U_{12}=D\prod_{(m,n)\in S}T_{m,n}(\theta_{m,n},\varphi_{m,n}),$ (12)
where $D$ represents diagonal phase factors at the output ports, $S$ denotes the ordered sequence of mode pairs, and each $T_{m,n}$ is the $12\times 12$ matrix with the $2\times 2$ block from Eq. (10) acting on modes $m$ and $n$. Applying this decomposition to our embedded evolution operators yielded the set of phase values $\{(\theta_{m,n},\varphi_{m,n})\}$ that parametrize each MZI on the chip.
Realizing these target phase shifts from the unitary decomposition requires compensating for phase variations in each MZI through systematic calibration. The full characterization and fitting procedure for the current-phase relation is provided in the Supplementary Materials (Sec. S2).
@@ 2.4. Data Acquisition and Analysis
We measured the transient population dynamics by applying the unitaries incrementally rather than as single operations. This allowed us to resolve intermediate dynamics throughout the modulation cycle. Each period $T$ was divided into 60 steps where 20 steps were allocated to each subprocess ($U_{1}$, $U_{2}$, $U_{3}$).
The measurement procedure began with state preparation using the input fiber switch set to the first port of the mesh. For the three-site and seven-site lattices, we initialized the system with unit population at site 1, implementing the state $|1\rangle$. The four-site system required the symmetric superposition $(|1\rangle+|3\rangle)/\sqrt{2}$, which we prepared using two sequential MZIs to route equal portions of the input light to modes 1 and 3.
For each discretized time step, we decomposed the corresponding unitary operator and applied the resulting phase shifts to the mesh. The current drivers set the required heater currents, with a 1-second dwell time ensuring thermal steady state before measurement. The initial state remained fixed throughout the evolution, with each unitary step representing the cumulative evolution from $t=0$ to the current time.
Output power measurements proceeded systematically across all active modes. The fiber switch automatically cycled through the output ports corresponding to the lattice sites, with a 50 ms switching delay ensuring mechanical stability. The optical power meter recorded the transmitted power at each port with hardware averaging of 20 samples. We treated each raw reading as having a $\pm 3\%$ relative uncertainty.
To account for variations in fiber-to-chip coupling efficiency, we normalized the measured powers at each time step. The population at site $i$ was calculated as:
$P_{i}=\frac{P_{i,\text{raw}}}{\sum_{j}P_{j,\text{raw}}},$ (13)
where $P_{i,\text{raw}}$ denotes the raw measured power at site $i$ and the sum extends over all active lattice sites. Assuming independent channel errors with $\sigma(P_{j,\text{raw}})=r\,P_{j,\text{raw}}$, the corresponding $1\sigma$ uncertainty on the normalized populations was obtained by standard error propagation:
$\sigma(P_{i})=r\,P_{i}\sqrt{\,1+\sum_{j}P_{j}^{2}-2P_{i}\,},\qquad r=0.03,$ (14)
and we report $P_{i}\pm\sigma(P_{i})$ throughout.

### v3

# arXiv:2509.13184v3[physics.optics] 03 Mar 2026

@@ Abstract
Photons, unlike electrons, do not couple directly to magnetic fields, yet synthetic gauge fields can impart magnetic-like responses and enable directional transport. Discretized Floquet evolution provides a controlled route, where the time-ordered sequencing of non-commuting Hamiltonians imprints complex hopping phases and breaks time-reversal symmetry. However, stabilizing such driven dynamics and observing unambiguous signatures of these effects on a reconfigurable platform has remained challenging. Here we demonstrate synthetic gauge fields for light on a programmable photonic processor by implementing discretized Floquet drives that combine static and dynamic phases. The resulting dynamics exhibit chiral circulation that reverses under drive inversion, flux-controlled interference with high visibility, and robust directional flow stabilized by optimizing the driving period. We further characterize the system using a first-harmonic phase as an order parameter, whose per-period winding number quantifies angular drift and reverses sign with the drive order. These results establish discretized Floquet evolution as a versatile framework for driven photonics, providing a programmable route to engineer gauge fields, stabilize driven phases, and probe winding-number signatures of chiral transport.
@@ 1. Introduction
Magnetic fields profoundly alter the behavior of charged particles by reshaping their energy spectrum. In two-dimensional electron systems, a perpendicular field condenses continuous bands into a hierarchy of discrete Landau levels, each separated by the cyclotron frequency and hosting a macroscopic degeneracy of states. This quantization underlies the quantum Hall effect, wherein the electrical conductance assumes precisely quantized values. The effect also produces chiral edge states, which are one-way electron channels along material boundaries that are topologically protected against backscattering from disorder [30, 14]. Photons propagate through such magnetic fields unchanged, their spectra unquantized, their transport unprotected. Remarkably, synthetic magnetic fields can induce analogous phenomena in photonic systems, thereby enabling magnetic-like responses and robust optical transport [32, 9, 12].
@@
In this work, we demonstrate a photonic simulator of synthetic magnetic fields based on discrete-time evolution. The simulator is implemented on a reconfigurable Mach-Zehnder interferometer (MZI) mesh, which supports arbitrary linear optical transformations [26, 5] and has enabled applications in quantum information processing [4, 23, 21, 33], neuromorphic computing [29, 28], and signal processing [22, 19]. We use a Floquet drive in which each period is divided into unitary substeps that sequentially activate selected waveguide pairs on the MZI mesh while other couplings are held off. The time-ordered product of these steps induces complex hopping phases for photons, realizing a synthetic magnetic flux. Because the substeps do not commute, the evolution depends on their order, which breaks time-reversal symmetry. This programmability allows a single photonic processor to investigate various driven phenomena without separate fabrications by simply adjusting control signals [15, 3], building on recent implementations of programmable lattices in integrated photonic meshes [20, 6, 18, 17]. Leveraging this versatility, we implement three-, four-, and seven-site lattice configurations. Our measurements reveal how each geometry produces distinct magnetic signatures, with the three-site system establishing the fundamental mechanism, the four-site system revealing flux interference, and the seven-site system demonstrating directional photon transport in complex systems.
@@ 2.1. Theoretical framework
Our approach utilizes discretized Floquet dynamics to induce synthetic gauge fields for photons. By subjecting the lattice to a time-ordered sequence of non-commuting unitary operations, we break time-reversal symmetry and imprint a path-dependent geometric phase on the photon transport.
Consider a programmable lattice governed by a periodic modulation cycle $T$. We partition the evolution into a sequence of discrete unitary operations $U_{k}$ (Fig. 1a,b), where each step activates a specific subset of couplings for a duration $T/3$ (Fig. 1c). While the instantaneous Hamiltonian at any step could be reciprocal, the non-commutativity of the sequence ($U_{k},U_{k^{\prime}}]\neq 0$) renders the total Floquet operator $U_{F}(T)$ distinct from its time-reversed counterpart. Physically, this temporal ordering allows us to define distinct clockwise (CW) and counter-clockwise (CCW) evolution operators (see section S1).
@@
Fig. 1: Discretized Floquet evolution on a programmable photonic processor. A state $|\psi(0)\rangle$ injected on the left bus enters a programmable photonic processor consisting of a reconfigurable MZI mesh (black box) that implements the time-evolution operator $U(t)$ on NN spatial waveguide modes, yielding $|\psi(t)\rangle=U(t)|\psi(0)\rangle$ at the right bus. a, CW drive: one period $T$ is decomposed into three sequential substeps $U_{1}(T/3)$ (red), $U_{2}(T/3)$ (green), and $U_{3}(T/3)$ (blue). These time-ordered operations constitute the Floquet operator $U_{F}^{\text{CW}}(T)=U_{3}U_{2}U_{1}$. Dashed guide lines indicate that this schedule is programmed onto the chip; horizontal dots indicate repetition over successive periods. At intermediate times (e.g., $t=T/3$), only the substep executed up to that moment has acted. b, CCW drive obtained by reversing the order of the substeps, giving $U_{F}^{\text{CCW}}(T)=U_{1}U_{2}U_{3}$. c, Scalable NN-site lattice representation. Lattice sites (white circles) map to physical waveguides, while solid colored bonds represent the tunable couplings realized by the MZI mesh active in the respective substeps (colors match panels a–b); dashed bond extensions indicate tiling to larger lattices. Curved arrows highlight the induced chiral circulation in each triangular plaquette, illustrating how the programmed drive order generates a synthetic magnetic flux.
To capture the complete transient dynamics of this driven system, we define the total time-evolution operator $U(t)$. For an arbitrary time $t=nT+\delta t$ (where $n$ is the number of completed cycles and $\delta t$ is the remainder, $0\leq\delta t<T$), the evolution is given by:
$U(t)=U_{\text{partial}}(\delta t)[U_{F}(T)]^{n},$ (1)
where $U_{F}(T)]^{n}$ describe the stroboscopic dynamics over full cycles, and $U_{\text{partial}}(\delta t)$ accounts for the residual intra-period evolution.
Stroboscopically, the system behaves as if governed by a time-independent effective Hamiltonian with complex couplings. The time-ordered drive ensures that a photon hopping between sites acquires a geometric phase analogous to the Aharonov-Bohm phase. The accumulation of these phases (e.g. around a closed triangular plaquette) generates a synthetic magnetic flux $\Phi_{\text{mod}}$. This flux relates formally to a synthetic vector potential $\mathbf{A}$, where the phase $\phi$ acquired along a link is given by the line integral $\phi=\int\mathbf{A}\cdot d\mathbf{l}$. Furthermore, we can augment this geometric phase by imprinting programmable static phases $\Phi_{\text{static}}$ onto the lattice bonds, allowing the total accumulated flux $\Phi_{\text{total}}=\Phi_{\text{mod}}+\Phi_{\text{static}}$ to be fully tunable (see section S1).
@@ 2.2. Experimental platform
The experimental platform (Fig. 2) was based on a 12-mode photonic processor. Here, the 12 modes represent independent spatial waveguide paths that constitute the Hilbert space of the system, where each waveguide acts as a discrete site in our programmed synthetic geometry. The processor integrates 264 active elements, consisting of a mesh of thermally tuned MZIs and multimode interferometers (MMIs). Deep trenches were included around the MZIs to suppress thermal crosstalk between neighboring phase shifters.
A pulsed diode laser driver provided the coherent light source at 1550 nm. Light was routed to individual chip inputs via a MEMS fiber optical switch optimized for TE polarization. Output signals were collected through a second identical switch connected to a high-sensitivity optical power meter with a germanium photodiode sensor. A PC coordinated all instruments to provide full automation. Phase relationships were preserved by stabilizing the chip temperature using a thermoelectric cooler and PID controller. Detailed technical specifications of the experimental platform are provided in sections S2 to S3.
2.3. Lattice realization and driving protocols
The programmable photonic chip allowed us to implement three distinct lattice configurations. We first realized a three-site triangular lattice whose single closed loop allowed clear interpretation of directional effects when driving CW and CCW, isolating the non-commuting operations that break time-reversal symmetry. We then programmed a four-site lattice consisting of two coupled triangular lattices with independent phase control. The shared vertices where pathways converge provided observation points for interference effects. Finally, we configured a seven-site hexagonal lattice to test synthetic magnetic fields in multi-loop geometries. Six perimeter sites surrounding a central node formed a network of interconnected triangular plaquettes, revealing whether directional transport persists in complex systems.
Operating conditions varied across lattice configurations based on their specific requirements. For coherent population transfer in the three-site and four-site systems, we set coupling strengths and evolution times to achieve Rabi $\pi$-pulses. This operating point is robust to perturbations in a broad range, given by a finite gap in the corresponding spectrum of the three-site Floquet operator, so the dynamics are tolerant to small period errors as elucidated in the Results (Section 3.3). While the triangular lattice experiments tested both modulation sequences to establish chirality, the four-site system used only CW driving to focus solely on the interference effects. The seven-site system required a different regime where Rabi pulses are not applicable, so we selected the modulation period by maximizing the minimal quasi-energy gap of the Floquet operator, which yields a stable evolution and robust directional flow. As in the 3-site case, testing both modulation sequences would confirm the presence of the synthetic magnetic field signatures.
@@ 2.4. Unitary implementation
We realized the evolution operators by decomposing the target unitary $U(t)$ into a sequence of local operations on the mesh. Each MZI implements a SU(2) transformation, parameterized by two independent phase degrees of freedom. To map the theoretical models onto the device, we defined the $n$ lattice sites as a subspace of the available spatial modes. The target $n\times n$ operator was embedded into the processor’s full Hilbert space by decoupling the inactive waveguides, ensuring the evolution remained confined to the synthetic lattice.
We then employed the Clements decomposition [5] to factorize this high-dimensional unitary into a deterministic arrangement of local phase shifts for the mesh. To ensure high-fidelity control, these target phases were mapped to heater currents using a calibrated nonlinear response model that compensates for fabrication variances. Further details regarding the unitary implementation are found in sections S4 to S6.
@@ 2.5. Data acquisition and analysis
To resolve the transient population dynamics, we applied the unitary evolution operators incrementally. Each modulation period $T$ was divided into 60 discretized steps (20 per subprocess). For a given time step $t$, the cumulative operator $U(t)$ was decomposed into phase configurations applied to the MZI mesh. The system was initialized in state $|1\rangle$ for the three-site and seven-site lattices, while the four-site lattice utilized a symmetric superposition $(|1\rangle+|3\rangle)/\sqrt{2}$ prepared via the first layers of the mesh. At each step, output powers were measured across all active waveguides using an optical power meter. To account for fiber-to-chip coupling variations, the raw power readings were normalized to the total detected power at that time step. We estimated experimental uncertainties by propagating a $\pm 3\%$ relative uncertainty through the normalization procedure (see section S7).